# Multiplicative and additive systematics in galaxy density fluctuations and clustering measurements

### v1

@@ 5.3 Multiple Systematics
We now move to the treatment of multiple systematics. When dealing with real survey data (e.g., DES), a multitude of systematic templates (Rodríguez-Monroy et al., 2022) are considered during the mitigation process, motivating our extention to the case of multiple systematics. We use a non-trivial number of uncorrelated contaminants, 25, to test our method. We generate 5 independent templates from each of the 5 systematic families described in Sec. 4.2, resulting in 25 independent templates. The contamination model used is chosen with a random number generator such that each systematic has equal chance of being additive or multiplicative, resulting in 10 additive and 15 multiplicative systematics for this analysis. The nonzero contamination parameters are chosen from a Gaussian centered at zero and standard deviation of $0.15$. This particular value of the standard deviation is motivated by real observations of seeing in HSC. Using these choices we produce the observed galaxy overdensity field for each realization of the mocks using equation (12). We then fit for the contamination parameters jointly for all systematics using the combined, additive, and multiplicative model, as well as the true contamination model for comparison. To clarify, in the additive and multiplicative case all systematics are assumed to follow that specific model, while in the true contamination model, each systematic follows their true assigned model when fitting.
We compare performance of the different likelihood models in Fig. 5. We show only the results using the Gaussian likelihood for the fits, as we found negligible difference when using the skewed Gaussian form. Both panels show that the combined model produces sub-percent mitigation of the contaminated two-point function (to a very small fraction of the statistical uncertainty) and equivalent performance to using the true contamination model within the statistical uncertainties. On the other hand, using the additive and multiplicative models consistently across all systematics results in biased corrections of the clustering signal (model mis-specification bias). We conclude, therefore, that assuming a single contamination model in the presence of systematics with distinct types of contamination can lead to biased estimates of the correlation function. In this case, the safest choice when the form of systematics is not known is to use the combined model for all of them.
@@
Figure 8: Top: Average parameter estimation variance for each systematics class used to generate contaminating templates. The variance for each estimated parameter is calculated over all 119 realizations of the KiDS-HOD mocks for each class of systematic. Since we use 5 systematics per class, the points show the average noise from the 5 parameters corresponding to each systematic contaminant. The different colors show the likelihood form used, either a Gaussian (blue) or a Skewed Gaussian (orange). The errorbars are the errors on the mean. The left panel shows the variance on the $a$ parameters (which modify the mean of the observed $\delta_{g}$ distribution), while the right panel shows the variance for the $b$ parameters (which modify its variance). As we can see, there is little difference between the estimated noise using either likelihood on the $a$ parameters. However, the Skewed Gaussian has lower parameter noise for the $b$ parameters. We attribute this to the Skewed Gaussian’s better recovery of the distribution’s variance when we do not remove outliers. Bottom: The effect of miscalculating the parameter noise on the two-point function correction. Possible parameters are drawn from a Gaussian centered at the true parameter value and with variance represented by the different colored lines. The correction to the observed correlation function is done using the estimated noise. The x-axis shows the ratio of the estimated noise to the true noise for both the $a$ and $b$ parameters, while the y-axis shows the bias in the corrected two-point function. We see that under- (or over-) estimating the parameter noise causes the two point function to be over- (or under-) corrected. However, the effect of the bias depends on the size of the noise itself. Having high parameter noise means the analysis is more sensitive to accurately estimating the noise, while the opposite is true at low parameter noise.
Contamination Model $\Delta\bar{A}\;(\times 10^{-2})$ $\sigma_{\bar{A}}\;(\times 10^{-2})$ Uncorrected $4.0$ $0.0135$ True $0.0016$ $0.0856$ Combined $0.0124$ $0.1188$ Multiplicative $-0.3974$ $0.1042$ Additive $1.4869$ $0.0367$
@@ 5.4 Cosmological Impact
To quantify the potential cosmological impact of how our method performs with multiple systematics, we fit the uncorrected and corrected clustering signals (for different modeling choices) to a constant times the true signal, as expressed in equation (25). Following the prescription detailed in Sec. 4.4, we calculate the estimated bias, $\Delta\bar{A}$, for different systematics contamination models in the multiple systematics case described above and report the results in Table 3. Although not shown, the differences between the estimated biases from using a Gaussian or skewed Gaussian likelihood are negligible, leading to a preference for the simpler Gaussian model. The bias-variance trade-off can be evaluated by comparing both columns in this table. As shown, the contaminated signal exhibits a few-percent bias that is highly statistically significant. Applying a correction using the true and the combined contamination models both reduce the bias to well below $0.1\sigma$, with the caveat that the error on the mean from the combined model is approximately 1.4 times larger than those of the true model (increased variance). As expected, using the true contamination model is the optimal approach when it comes to reducing the bias without much increase in the variance. However, the combined model still dramatically outperforms the other contamination models and produces a bias consistent with zero, indicating that it is a sufficient correction for a mixture of different types of systematic contaminants for cosmological analysis.
@@ 6.1 Comparison with other methods
The method presented in this paper also considers distinct parameters contributing to the additive and multiplicative contamination to the galaxy overdensity, allowing for empirical determination of the systematics model, rather than fitting a specific model (e.g., additive or multiplicative). This not only allows for the possibility of multiple sources of contamination for a given template, but also allows us to directly determine the functional form of the systematic contamination. The contamination parameters are also fitted for jointly, rather than independently for every systematic. This differs from previous iterative methods, which clean contaminants one by one using a form of linear regression. These methods rely on null hypothesis testing during the iterative process to avoid biased mitigation of contaminants. In contrast, our method deals with this problem by taking into account all systematics simultaneously in the fitting process.
@@
Finally, our method allows for fitting the $\delta_{g}$ distribution using a Skewed Gaussian likelihood, which more closely resembles the real distribution. Even though we do not see any significant difference in systematics mitigation between methods after debiasing the parameters, we still provide an alternative fit to a distribution commonly assumed to be Gaussian for systematics treatment. We emphasize the non-Gaussianity of the $\delta_{g}$ distribution in the fits for the true distribution in Sec. 5.1. There, the parameter that describes skewedness was estimated to be $\gamma\approx 5.1\pm 0.2$, which represents a highly non-Gaussian distribution. Despite the high statistical significance with which the non-Gaussianity of the $\delta_{g}$ distribution is detected, the effectiveness of systematics mitigation for the Gaussian likelihood indicates that for our scenario, model mis-specification of a non-Gaussian distribution as a Gaussian is not a problem when parameter noise is known or can be estimated accurately.
@@ 6.2 Practical applications on real data
Before applying data-driven methods to estimate the noise in the estimated systematics parameters, it is important to understand exactly what the noise depends on and how to evaluate its impact on systematics mitigation. The noise can depend on many factors, including the sky area coverage, the $\delta_{t}$ distributions and power spectrum, and the form of the likelihood used to analyze the $\hat{\delta}_{g}$ distribution (including the number of parameters being fitted). For example, the top panel of Fig. 8 illustrates how the noise level in the inferred systematics parameters can vary based on the form of the likelihood (blue and orange points represent results with a Gaussian and skewed Gaussian likelihood, respectively) and on the spatial power spectra used to generate the systematic maps, normalized to a fixed variance in $\delta_{t}$. As we can see, the noise can differ greatly depending on the power spectrum used, with some forms increasing the estimated noise variance by almost an order of magnitude. The choice of a Gaussian or skewed Gaussian likelihood does not significantly affect the noise in the $a$ parameters, which control shifts in the mean of the distribution. However, the choice has a substantial impact on the noise in the estimated $b$ parameters, which modulate its variance. Therefore, even though the choice of the likelihood (Gaussian or skewed Gaussian) does not significantly affect the systematics mitigation outcome, it is important to note that the noise estimates differ. However, we found no correlation between the noise variance in the estimated systematics parameter and the true value of the parameter.
[…]
@@
[…]
@@
Note that we cannot simply take the parameter noise variance directly from a single realization by looking at the posterior generated from the MCMC samples. Because of our simple adopted likelihood model, which neglects correlations in overdensity between pixels, the variance of the posterior coming from the MCMC samples does not reflect sources of uncertainty such as cosmic variance. Using the MCMC sample posterior noise will produce underestimates of the parameter noise and therefore lead to overcorrection of the two-point function.
@@ 7 Conclusions
In this paper, we presented a new method for characterizing and mitigating systematic contaminants in in galaxy overdensity fields. We introduced a generalized and flexible contamination model that allows for multiple systematics to contaminate the overdensity field in additive, multiplicative, or ‘combined’ forms. This modelling choice allows us to quantify contamination without assuming that all systematics follow one particular form, and in addition helps us empirically determine the functional form of systematics. Using the generalized model can reduce model biases, producing unbiased estimates of the two-point correlation function in the cases where different systematics take different forms captured by our model. The ability to determine the functional form of the contamination may provide insight into their origin, so they can be reduced or entirely avoided in future.
@@
[…]
@@ Appendix A Correlated Systematics
$w_{\text{corr}}=\frac{\hat{w}-\sum_{i}^{N_{\text{sys}}}a_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}{1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{2^{\prime}}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}$ (31)
@@
$\hat{w}(\theta)$ $=w(\theta)\left(1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle+2\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N_{\text{sys}}}b_{i}b_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle\right)$
@@
$w_{\text{corr}}=\frac{\hat{w}-\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle-\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N}a_{i}a_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle}{1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle+2\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N_{\text{sys}}}b_{i}b_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle}$ (32)

### v2

@@ 5.3 Multiple Systematics
We now move to the treatment of multiple systematics. When dealing with real survey data (e.g., DES), a multitude of systematic templates (Rodríguez-Monroy et al., 2022) are considered during the mitigation process, motivating our extention to the case of multiple systematics. We use a non-trivial number of uncorrelated contaminants, 25, to test our method. We generate 5 independent templates from each of the 5 systematic families described in Sec. 4.2, resulting in 25 independent templates. The contamination model used is chosen with a random number generator such that each systematic has equal chance of being additive or multiplicative, resulting in 10 additive and 15 multiplicative systematics for this analysis. The nonzero contamination parameters are chosen from a Gaussian centered at zero and standard deviation of $0.15$. This particular value of the standard deviation is motivated by real observations of seeing in HSC and chosen to result in an approximate 10% contamination level at larger scales when using the systematic families described in Subsection 4.2. This level of contamination is reasonable considering the results of Rodríguez-Monroy et al. (2022). Using these choices we produce the observed galaxy overdensity field for each realization of the mocks using equation (12). We then fit for the contamination parameters jointly for all systematics using the combined, additive, and multiplicative model, as well as the true contamination model for comparison. To clarify, in the additive and multiplicative case all systematics are assumed to follow that specific model, while in the true contamination model, each systematic follows their true assigned model when fitting.
We compare performance of the different contamination models in Fig. 5. We show only the results using the Gaussian likelihood for the fits, as we found negligible difference when using the skewed Gaussian form. Both panels show that the combined model produces sub-percent mitigation of the contaminated two-point function (to a very small fraction of the statistical uncertainty) and equivalent performance to using the true contamination model within the statistical uncertainties. On the other hand, using the additive and multiplicative models consistently across all systematics results in biased corrections of the clustering signal (model mis-specification bias). We conclude, therefore, that assuming a single contamination model in the presence of systematics with distinct types of contamination can lead to biased estimates of the correlation function. In this case, the safest choice when the form of systematics is not known is to use the combined model for all of them. We note that the increased degradation at larger scales is due to the small area of the KiDS mocks. This was confirmed using larger simulated mocks described Sec. 6.2, where this degradation is not present at these scales. We also verified that the combined model can still recover the truth and outperforms the simpler models in the case of no contamination ($a=b=0$) for a random selection of templates.
@@
Figure 8: Top: Average parameter estimation variance for each systematics class used to generate contaminating templates. The variance for each estimated parameter is calculated over all 119 realizations of the KiDS-HOD mocks for each class of systematic. Since we use 5 systematics per class, the points show the average noise from the 5 parameters corresponding to each systematic contaminant from the mocks. The crosses show the parameter variance taken from the MCMC posteriors for comparison. The different colors show the likelihood form used, either a Gaussian (blue) or a Skewed Gaussian (orange). The errorbars are the errors on the mean. The left panel shows the variance on the $a$ parameters (which modify the mean of the observed $\delta_{g}$ distribution), while the right panel shows the variance for the $b$ parameters (which modify its variance). As we can see, there is little difference between the estimated noise using either likelihood on the $a$ parameters. However, the Skewed Gaussian has lower parameter noise for the $b$ parameters. We attribute this to the Skewed Gaussian’s better recovery of the distribution’s variance when we do not remove outliers. In addition, we see that the variance recovered from the MCMC posteriors vastly underestimates the parameter noise. Bottom: The effect of miscalculating the parameter noise on the two-point function correction. Possible parameters are drawn from a Gaussian centered at the true parameter value and with variance represented by the different colored lines. The correction to the observed correlation function is done using the estimated noise. The x-axis shows the ratio of the estimated noise to the true noise for both the $a$ and $b$ parameters, while the y-axis shows the bias in the corrected two-point function. We see that under- (or over-) estimating the parameter noise causes the two point function to be over- (or under-) corrected. However, the effect of the bias depends on the size of the noise itself. Having high parameter noise means the analysis is more sensitive to accurately estimating the noise, while the opposite is true at low parameter noise.
Contamination Model $\Delta\bar{A}\;(\times 10^{-2})$ $\sigma_{\bar{A}}\;(\times 10^{-2})$ Uncorrected $4.0$ $0.0135$ True -0.0123 0.0631 Combined -0.0166 0.0861 Multiplicative 0.1123 0.0728 Additive 0.7214 0.0374
@@ 5.4 Cosmological Impact
To quantify the potential cosmological impact of how our method performs with multiple systematics, we fit the uncorrected and corrected clustering signals (for different modeling choices) to a constant times the true signal, as expressed in equation (25). Following the prescription detailed in Sec. 4.4, we calculate the estimated bias, $\Delta\bar{A}$, for different systematics contamination models in the multiple systematics case described above and report the results in Table 3. Although not shown, the differences between the estimated biases from using a Gaussian or skewed Gaussian likelihood are negligible, leading to a preference for the simpler Gaussian model. The bias-variance trade-off can be evaluated by comparing both columns in this table. As shown, the contaminated signal exhibits a few-percent bias that is highly statistically significant. Applying a correction using the true and the combined contamination models both reduce the bias to well below $0.1\sigma$, with the caveat that the error on the mean from the combined model is approximately 1.4 times larger than that of the true model (increased variance). As expected, using the true contamination model is the optimal approach when it comes to reducing the bias without much increase in the variance. However, the combined model still dramatically outperforms the other contamination models and produces a bias consistent with zero, indicating that it is a sufficient correction for a mixture of different types of systematic contaminants for cosmological analysis.
@@ 6.1 Comparison with other methods
The method presented in this paper also considers distinct parameters contributing to the additive and multiplicative contamination to the galaxy overdensity, allowing for empirical determination of the systematics model, rather than fitting a specific model (e.g., additive or multiplicative). This not only allows for the possibility of multiple sources of contamination for a given template, but also allows us to directly determine the functional form of the systematic contamination. The contamination parameters are also fitted for jointly, rather than independently for every systematic. This differs from iterative methods, which clean contaminants one by one using a form of linear regression. These methods rely on null hypothesis testing during the iterative process to avoid biased mitigation of contaminants. Other methods like ElasticNet incorporate the impact of multiplicative effects by using a two-step process to update the covariance for the assumed multiplicative parameters. In contrast, our method deals with this problem by taking into account all systematics simultaneously in the fitting process and self-consistently incorporating the multiplicative effects into the likelihood. This joint estimation enables the discrimination of additive and multiplicative templates directly.
@@
Finally, our method allows for fitting the $\delta_{g}$ distribution using a Skewed Gaussian likelihood, which more closely resembles the real distribution. Even though we do not see any significant difference in systematics mitigation between methods after debiasing the parameters, we still provide an alternative fit to a distribution commonly assumed to be Gaussian for systematics treatment(see examples of non-Gaussinity in Rezaie et al., 2021, 2023). We emphasize the non-Gaussianity of the $\delta_{g}$ distribution in the fits for the true distribution in Sec. 5.1. There, the parameter that describes skewness was estimated to be $\gamma\approx 5.1\pm 0.2$, which represents a highly non-Gaussian distribution. We note that the specific level of non-Gaussianity will heavily depend on the pixelazion scheme and galaxy density. Despite the high statistical significance with which the non-Gaussianity of the $\delta_{g}$ distribution is detected, the effectiveness of systematics mitigation for the Gaussian likelihood indicates that for our scenario, model mis-specification of a non-Gaussian distribution as a Gaussian is not a problem when parameter noise is known or can be estimated accurately.
@@ 6.2 Practical applications on real data
Before applying data-driven methods to estimate the noise in the estimated systematics parameters, it is important to understand exactly what the noise depends on and how to evaluate its impact on systematics mitigation. The noise can depend on many factors, including the sky area coverage, the $\delta_{t}$ distributions and power spectrum, and the form of the likelihood used to analyze the $\hat{\delta}_{g}$ distribution (including the number of parameters being fitted). For example, the top panel of Fig. 8 illustrates how the noise level in the inferred systematics parameters can vary based on the form of the likelihood (blue and orange points represent results with a Gaussian and skewed Gaussian likelihood, respectively) and on the spatial power spectra used to generate the systematic maps, normalized to a fixed variance in $\delta_{t}$. As we can see, the noise can differ greatly depending on the power spectrum used, with some forms increasing the estimated noise variance by almost an order of magnitude. The choice of a Gaussian or skewed Gaussian likelihood does not significantly affect the noise in the $a$ parameters, which control shifts in the mean of the distribution. However, the choice has a substantial impact on the noise in the estimated $b$ parameters, which modulate its variance. Therefore, even though the choice of the likelihood (Gaussian or skewed Gaussian) does not significantly affect the systematics mitigation outcome, it is important to note that the noise estimates differ. This could be relevant in the scenario where the parameter variance cannot be accurately measured and corrected for, in which case it is helpful to adopt a model with smaller variance since the bias correction would be less sensitive to it. However, we found no correlation between the noise variance in the estimated systematics parameter and the true value of the parameter.
In the absence of mock catalogs, we propose the use of the block bootstrap method to estimate the parameter noise. Block bootstrap means that the pixelated maps are divided into approximately equal area patches, where each bootstrap sample represents a version of the maps made from a random selection (with replacement) of the patches. This allows us to create multiple realizations from a single map and estimate the parameter noise one would get from having multiple independent mocks. We wish to test this data-driven approach on higher and more realistic area coverage than our current KiDS mocks. To do so we use the same mechanism to generate systematic template maps from the set of 5 systematic classes described in the multiple systematic case in Sec. 5.3. In order to obtain different area coverage $\delta_{g}$ maps, we used version 2.3.0 of CCL [^6] https://github.com/LSSTDESC/CCL (Chisari et al., 2019) to generate galaxy-galaxy power spectra using a vanilla $\Lambda$CDM cosmological model ($\Omega_{m}=0.3,\Omega_{b}=0.05,\;\Omega_{\Lambda}=0.7,\;\sigma_{8}=0.8,\;n_{s}=0.96$) with galaxy bias equal to $1$ ($\delta_{m}=\delta_{g}$) to produce Gaussian galaxy overdensity maps. This cosmological model is similar (while not identical) to that used in KiDS mocks, with the exception that our galaxy bias is fixed to 1. We use a Gaussian $n(z)$ with mean redshift $\mu_{z}=0.3$ and $\sigma_{z}=0.05$ to resemble the redshift range used for the previous validation with the KiDS mocks. We choose 4 size configurations: 100 deg2, 400 deg2, 1000 deg2, and 2000 deg2, to show how this approach works at various sky area coverages. Although not identical, the produced mocks provide a realistic test of the block bootstrap method for estimating parameter noise. We wish to compare the parameter noise estimation using this bootstrap approach with estimates using multiple realizations. We transformed the Gaussian maps into log-normal maps to more accurately portray a realistic underlying distribution. This is done by taking some Gaussian map $G$, shifting it using its variance: $S=G-0.5*\text{Var}[G]$, and transforming the shifted map to a log-normally distributed one: $L=\exp(S)-1$. For the tests described here the shift parameter is $0.0250.5*\text{Var}[G]=$ 0.025.
@@
In Fig. 9 we compare the accuracy of parameter noise estimation by comparing a bootstrap approach on a single mock catalog realization against our previous approach using multiple realizations of mock catalogs as a function of sky area coverage. As we can see from the figure, the noise is approximately proportional to the inverse of the survey area. The figure also shows that using the block bootstrap on a single dataset can be a good alternative to using multiple realizations of mock catalogs. The bootstrap estimates are quite good for higher area mocks, and for small area mocks with lower levels of parameter noise. However, the approach fails for small area surveys ($\sim$100 deg2) with high parameter noise associated with systematics maps that have a lot of small-scale power.
@@
Note that we cannot simply take the parameter noise variance directly from a single realization by looking at the posterior generated from the MCMC samples. Because of our simple adopted likelihood model, which neglects correlations in overdensity between pixels, the variance of the posterior coming from the MCMC samples does not reflect sources of uncertainty such as cosmic variance. Using the MCMC sample posterior noise will produce underestimates of the parameter noise, as seen in Fig. 8, and therefore lead to overcorrection of the two-point function.
@@ 7 Conclusions
In this paper, we presented a new method for characterizing and mitigating systematic contaminants in in galaxy overdensity fields. We used a generalized and flexible contamination model that allows for multiple systematics to contaminate the overdensity field in additive, multiplicative, or ‘combined’ forms. This modelling choice allows us to quantify contamination without assuming that all systematics follow one particular form, and in addition helps us empirically determine the functional dependence of systematics on the galaxy overdensity. Using the generalized model can reduce model biases, producing unbiased estimates of the two-point correlation function in the cases where different systematics take different forms captured by our model. The ability to determine the functional form of the contamination may provide insight into their origin, so they can be reduced or entirely avoided in future.
@@
Finally, we considered the practical applications of the methods presented here in real data. The main challenge is to accurately estimate the systematics model parameter noise needed to obtain unbiased corrected clustering signals. Since this work used mock catalogs, parameter noise could be estimated well from multiple mock catalog realizations. However, this will pose a challenge when the method is applied to real data, as mock catalogs are not always available and are computationally expensive. To overcome this challenge, we proposed a block bootstrap approach for estimating the parameter noise from a single realization of both the galaxy density field and systematic maps. We showed that the estimated parameter noise using the block bootstrap is comparable to the estimates using multiple mock catalog realizations in all cases except for small area ($\sim$100 deg2) with significant small-scale power in the systematics template maps. This is therefore a promising approach in applying this method to real data in future work.
@@ Appendix A Correlated Systematics
$w_{\text{corr}}=\frac{\hat{w}-\sum_{i}^{N_{\text{sys}}}a_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}{1+\sum_{i}^{N_{\text{sys}}}b_{i}^{2^{\prime}}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}$ (31)
@@
$\hat{w}(\theta)$ $=w(\theta)\left(1+\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle+\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N_{\text{sys}}}b_{i}b_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle\right)$
@@
$w_{\text{corr}}=\frac{\hat{w}-\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle-\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N}a_{i}a_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle}{1+\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle+\sum_{i}^{N_{\text{sys}}}\sum_{j\neq i}^{N_{\text{sys}}}b_{i}b_{j}\langle\delta_{t_{i}}\delta_{t_{j}}\rangle}$ (32)